Length of a string of density rho and Young’s modulus y under tension is increased by 1/n times of its original length. If the velocity of transwerse and longitudinal vibrations of string is same, find the value of such velocity?

Grade:11

1 Answers

Khimraj
7 years ago
y = Tl/Adl and dl = l/n so y = Tn/A and y = (m/Al)n[T/(m/l)] then y = ρnT/u and as T/u is velocity. So y= ρnv. Then velocity v = y/ρn.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer